An Alkene on ozonolysis gives a mixture of formaldehyde and acetone. predict the structural formula of that alkene. Give the IUPAC name.


 


Please reply.


Thanks.

Grade:

4 Answers

Rituraj Verma
12 years ago


CH3-C(CH3)=CH2   IUPAC NAME: 2-methyl propene

 

mohammed saquib iqbal
12 years ago

iupac name= i-phenylprop-1-ene

Gitika
6 years ago
2 Methyl prop-1-ene is the iupac name of that alkene and it`s structural formula is ch3-c(ch3)=ch2.this process is known as ozonolysis
Kushagra Madhukar
3 years ago
Dear student,
Please find the answer to your question.
 
2 Methyl prop-1-ene is the iupac name of the alkene.
It’s structural formula is CH3–C(CH3)=CH2.this process is known as ozonolysis
 
Thanks and regards,
Kushagra